Lumber Company produces lumber for many applications. The company had the following data for July:

Units

Total

Per Unit

Sales

14500

826500

57

Variable Expenses

493000

34

Contribution Margin

333500

23

Fixed Expenses

165000

Net Operating Income

168500

During a meeting with senior leadership, you are asked about several possible scenarios regarding possible changes in sales and costs. Use the spreadsheet to calculate all the scenarios to bring back to the next meeting.

Activity cost pool Activity Measure Activity Rate
Order size Direct labor hours $14 per direct labor hour
Customer orders Customer orders $114 per customer order
Quality testing Test hours $74 per hour
Sales Sales calls $634 per sales call
Order #1 required 475 direct labor hours, 9 hours of quality testing, and 7 sales calls
Order #2 required 115 direct labor hours, 48 hours of quality testing, and 19 sales calls
Total overhead cost assigned
Activity Cost Pool Order #1 Order #2
Total Overhead cost
Answer #1

Answer

Total overhead cost assigned
Activity Cost Pool Order #1 Order #2
Order size $6,650 $1,610
Customer Orders $114 $114
Quality Testing $666 $3,552
Sales $4,438 $12,046
Total Overhead cost $11,868 $17,322

--Working

Total overhead cost assigned
Activity Cost Pool Order #1 Order #2
Order size =475*14 =115*14
Customer Orders =1*114 =1*114
Quality Testing =9*74 =48*74
Sales =7*634 =19*634